# Deep-Learning Control of Lower-Limb Exoskeletons via simplified Therapist Input

### v1

@@ IV Conclusions
In this work, we present a three-step data-driven approach designed to replace the hierarchical structure of traditional exoskeleton controllers. The first step involves probabilistically mapping a short-term history of sensor data (including encoders, force sensors, and foot-plate sensors) to a minimal set of clinically relevant gait features: step length, height (to differentiate ascending and descending surfaces), walking velocity, step clearance, and gait phase. These features can either be applied directly or adjusted by an operator/therapist via a user interface to customize the user’s locomotion. The final step uses probabilistic regression to map the current locomotion features to the desired joint posture of the exoskeleton, leveraging prediction uncertainty to adapt the impedance of a spring-damper system linking the user and the device. This adaptive compliance enhances interaction safety. Validation on two healthy users across diverse gait patterns demonstrated the ability of the proposed approach of detecting the user intent and providing assistance accordingly, generating predominantly negative interaction power between the user and the exoskeleton. Future work will explore the selection of additional clinically relevant features in collaboration with therapists to ensure that they align with practical rehabilitation needs. Additionally, more extensive studies involving patient populations (e.g., stroke, spinal cord injury) are necessary to evaluate the impact of these assistive strategies on individuals with gait impairments.

### v2

@@ IV Conclusions
In this work, we present a three-step data-driven approach designed to replace the hierarchical structure of traditional exoskeleton controllers. The first step involves probabilistically mapping a short-term history of sensor data (including encoders, force sensors, and foot-plate sensors) to a minimal set of clinically relevant gait features: step length, height (to differentiate ascending and descending surfaces), walking velocity, step clearance, and gait phase. These features can either be applied directly or adjusted by an operator/therapist via a user interface to customize the user’s locomotion. The final step uses probabilistic regression to map the current locomotion features to the desired joint posture of the exoskeleton, leveraging prediction uncertainty to adapt the impedance of a spring-damper system linking the user and the device. This adaptive compliance enhances safety, limiting large desired interaction torques when the prediction of the user’s behavior is less accurate. Validation on two healthy users across diverse gait activities demonstrated the ability of the proposed approach to detect user intent and provide assistance accordingly, generating predominantly negative interaction power between the user and the exoskeleton. Future work will explore the selection of additional clinically relevant features in collaboration with therapists to ensure that they align with practical rehabilitation needs. Additionally, more extensive studies involving patient populations (e.g., stroke, spinal cord injury) are necessary to evaluate the impact of these assistive strategies on individuals with gait impairments.